OpenStudy (anonymous):

plz help gold medal will be awarded. if a chemical reaction catalyzed by an enzyme is being carried out, and there is a sudden, drastic decrease in temperature, what is most likely to happen next?

OpenStudy (shrutipande9):

See enzyme are temperature dependent..they work only in a particular temperature range.....so what do u think

OpenStudy (anonymous):

oh thanks so much

OpenStudy (anonymous):

oh um i dont know

OpenStudy (anonymous):

enzyme activity will be decreased bcz enzymes get denatured at low and hgh temperatures..

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

thx

OpenStudy (frostbite):

You wrote I should contribute, but I'm a physical biochemist so I give a mathematical answer: For a system that follow the Arrhenius equation we get that: \[\large k=A ~ \exp \left( \frac{ -E _{a} }{ k _{b} \times T } \right)\] What happens if T increases? Assume no denaturation.

OpenStudy (anonymous):

???

OpenStudy (anonymous):

omg frostbite....

OpenStudy (anonymous):

i haven't seen this equation before :D

OpenStudy (frostbite):

In other words: